# Efficient Model-Based Reinforcement Learning Through Optimistic Thompson Sampling

### v1

@@ Abstract
Learning complex robot behavior through interactions with the environment necessitates principled exploration. Effective strategies should prioritize exploring regions of the state-action space that maximize rewards, with optimistic exploration emerging as a promising direction aligned with this idea and enabling sample-efficient reinforcement learning. However, existing methods overlook a crucial aspect: the need for optimism to be informed by a belief connecting the reward and state. To address this, we propose a practical, theoretically grounded approach to optimistic exploration based on Thompson sampling. Our model structure is the first that allows for reasoning about joint uncertainty over transitions and rewards. We apply our method on a set of MuJoCo and VMAS continuous control tasks. Our experiments demonstrate that optimistic exploration significantly accelerates learning in environments with sparse rewards, action penalties, and difficult-to-explore regions. Furthermore, we provide insights into when optimism is beneficial and emphasize the critical role of model uncertainty in guiding exploration.
@@ 3 Problem Statement
for transition dynamics $f_{t}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and reward function $f_{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Our objective is to learn optimal control for this system within episodes of finite time horizon $T$. To control the agent, we learn a deterministic or stochastic policy $\pi$ that selects actions given the current state. For ease of notation, we write the policy deterministically as $\pi:\mathcal{S}\rightarrow\mathcal{A}$ so that $a_{t}=\pi(s_{t})$. We consider a specific transition dynamic $\hat{f}_{t}$ and reward function $\hat{f}_{r}$, denoted with a hat to indicate realizations that are drawn from random functions. The performance of the policy is the sum of accumulated rewards during an episode,
@@ 5.1 Model Structure
where $\mu_{s}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and $\mu_{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ are the dynamics and reward mean functions and $K_{ss},K_{sr},K_{rr}$ represent covariance functions $\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Traditional GPs describe outputs as conditionally independent given the input and hence factorize across the state dimensions and the reward. However, we hypothesize that this assumption is unfounded in most RL tasks and furthermore antithetical to principled optimistic exploration. Therefore, we want to explicitly model the correlation structure between the dimensions of the predicted state, and more importantly, between the predicted state and its associated reward. While this correlation can naturally be included in the GP by parameterizing the full covariance structure across examples and output dimensions, this is computationally intractable due to cubic scaling with the size of the covariance matrix (Rasmussen & Williams, 2006). Instead, we adopt a linear model of coregionalization (Grzebyk & Wackernagel, 1994; Wackernagel, 2003) where the covariance between output dimensions is restricted to be linear combinations of the input covariance. This induces a Kronecker structured covariance function that allows for efficient factorization while still meaningfully directing optimistic exploration.
In most GP regression use cases, the mean function is a constant function. While this is a valid modeling assumption for stationary data, the dynamics and reward structures we encounter in RL tasks vary substantially across different regions of the input space. A stationary assumption would lead to significant model mismatch and transfer the responsibility of the mean to the covariance function, causing the model to badly represent the uncertainty, resulting in inefficient exploration and misguided optimism. To circumvent this problem, we use a multi-layer MLP as a mean function as in Iwata & Ghahramani (2017); Fortuin et al. (2019). This allows learning a flexible mean function for non-stationary transitions while providing uncertainty that can be used in an optimistic fashion. To learn the covariance and mean functions, we perform approximate inference by optimizing a variational lower bound on the marginal likelihood, as done by Titsias (2009). This updates the posterior distribution $p(\hat{f}\mid\mathcal{D}_{\text{env}})$, which we use to obtain the posterior predictive distribution $p(s_{t+1},r_{t}\mid s_{t},a_{t})$. Given the current state $s_{t}$ and action $a_{t}$, this multivariate Gaussian distribution reflects the model’s predictions and uncertainty over the next state and reward.
@@ 5.2 Using Thompson Sampling in the GP Acquisition Function
Figure 1: Learning curves on MuJoCo tasks averaged over 5 seeds. HOT-GP demonstrates equivalent or superior sample efficiency and performance for all tasks considered. The dashed line denotes SAC performance at convergence within 1,000,000 environment steps (3,000,000 for Half-Cheetah).
@@ 6 Experiments
We evaluate the performance of our proposed algorithm, HOT-GP, on continuous state-action control tasks. Since inaccurate reward estimates are likely initially, we increase $r_{\text{min}}$ linearly throughout training to become more optimistic as the model becomes more reliable. In our experiments, we consider three variations of HOT-GP where $r_{\text{min}}$ begins at 0.1 and increases to either 0.3, 0.5, or 0.7. We also consider the alternative Thompson sampling and greedy exploration strategies. Our Thompson sampling implementation follows the HOT-GP procedure but draws samples from the reward-conditioned distribution $p(s_{t+1}\mid s_{t},a_{t},r_{t}=\hat{r}_{t})$ rather than taking the expectation. For greedy exploitation, we predict the next state and reward as the mean from the model.
First, we compare these approaches on standard MuJoCo benchmark tasks (Todorov et al., 2012) against additional model-free and model-based methods. Then we evaluate these approaches on a practical robotics task using the VMAS simulator (Bettini et al., 2022) and provide an analysis of the individual components. In-depth algorithmic descriptions of HOT-GP, along with the model-based methods we use for comparison, are given in Appendix A. Implementation details about our training procedures are discussed in Appendix B and specific tasks are described in Appendix C.
@@ 6.1 MuJoCo Benchmarks
Sparse Rewards. This task provides reward $r_{\text{dist}}(s)$ that is nonzero only when the agent is within a small threshold of the target location, thus demanding clever exploration strategies to reach the target as there is no dense feedback. As in Curi et al. (2020), we introduce an action penalty of the form $r(s,a)=r_{\text{dist}}(s)-\rho\cdot\text{cost}(a)$ with the aim of penalizing random and encouraging local exploration. This is described in greater detail in Appendix C. We evaluate HOT-GP and the other comparison methods on Sparse Reacher using three different action penalty weights: $\rho=0$, $\rho=0.1$, and $\rho=0.5$, and present the results in the bottom row of Figure 1. HOT-GP demonstrates significantly superior sample efficiency over MBPO and moderately improved sample efficiency over optimistic MBPO in each case. The findings on Sparse Reacher with $\rho=0$ indicate that all the HOT-GP variants perform robust optimistic exploration and successfully uncover the signal despite sparse rewards. When using the action penalty $\rho=0.1$, HOT-GP with $r_{\text{min}}$ from 0.1 to 0.3 performs poorly relative to higher levels of optimism. This suggests that higher levels of optimism are necessary to overcome the prior against taking large actions and explore the reward signal $r_{\text{dist}}$. HOT-GP is able to do this by relating the action-based penalty to the state to the overall reward through the joint model. The less efficient learning of optimistic MBPO highlights the critical importance of having a joint model of the reward and state distributions to facilitate efficient exploration. For $\rho=0.5$, we observe that the cost term typically accounts for nearly half of the overall reward. As a result, the HOT-GP variants explore much of the cost signal rather than $r_{\text{dist}}$, leading to their similar performance in this setting.
@@ 7 Conclusion
Future Work. Future work should study how the schedule on $r_{\text{min}}$ affects optimism and principled ways for setting this value throughout training, as we only considered a linear schedule. Another promising avenue would extend the planning horizon. We concentrated on learning 1-step lookahead reward-dynamics as HOT-GP can implicitly propagate uncertainty across the horizon by using pointwise uncertainty estimates from the model. Nevertheless, there may be benefits in explicitly propagating uncertainty to facilitate optimism over longer-term reward predictions (Seyde et al., 2021; Mehta et al., 2022; Hansen et al., 2024). Furthermore, our investigation indicates that the utility of optimistic exploration depends on the complexity of the task, although our experiments were limited to five tasks. Future research should explore the role of optimism across a broader range of problems, including in real-world scenarios with on-robot learning.
@@ Acknowledgments
J. Bayrooti is supported by a DeepMind scholarship. A. Prorok is supported in part by European Research Council (ERC) Project 949940 (gAIa). We also thank Aidan Scannell for insightful conversations and useful feedback on this work.
@@ Appendix A Algorithm Details
In this section, we provide more detailed information about the algorithms we use for evaluation.
@@ A.1 Model-Free Baselines
PPO. We also use a standard version of PPO as introduced in Schulman et al. (2017).
@@ Appendix B Training
GP Training. Our approach requires access to the full posterior distribution of the Gaussian process. To facilitate this in an efficient manner, we train the GP using a variational bound based on inducing points (Hensman et al., 2015) following the approach given in Iwata & Ghahramani (2017). This allows for efficient training using stochastic inference. To further reduce the computational cost of learning the model, we do not train on the whole dataset on each iteration and instead we draw $1,000$ randomly sampled transitions from $\mathcal{D}_{\text{env}}$. This approach reduces the training cost significantly and provides an additional source of stochasticity that helps training. Using a Gaussian likelihood allows us to compute the predictive posterior in closed form, allowing us direct access to its mean and variance.
Hyperparameters. We train all algorithms using a set of hyperparameters that we found to be optimal through trial-and-error or were recommended. Table 1 describes the hyperparameter values specific to the MuJoCo and VMAS tasks we considered. We also report additional algorithm-specific hyperparameters used. In H-UCRL, we use the exploration-exploitation coefficient $\beta=0.01$ and number of samples $Z=5$. In MBPO and optimistic MBPO, we use an ensemble of size 7. For DDPG, we use an exploration noise $\sigma_{\text{explore}}$ from 1 to 0.1. For GP-based approaches, we use the Matern kernel as the covariance function and learn from 100 inducing points. For the MuJoCo tasks, we use a 4-layer MLP with 200 hidden units per layer and SiLU activation function. For the VMAS coverage task we use a 2-layer MLP with 200 hidden units per layer and Mish activation function. We use the Adam optimizer in all cases.
@@
Steps per model rollout $K$ 1 1 1 1 1

### v2

@@ Abstract
Learning complex robot behavior through interactions with the environment necessitates principled exploration. Effective strategies should prioritize exploring regions of the state-action space that maximize rewards, with optimistic exploration emerging as a promising direction aligned with this idea and enabling sample-efficient reinforcement learning. However, existing methods overlook a crucial aspect: the need for optimism to be informed by a belief connecting the reward and state. To address this, we propose a practical, theoretically grounded approach to optimistic exploration based on Thompson sampling. Our approach is the first that allows for reasoning about joint uncertainty over transitions and rewards for optimistic exploration. We apply our method on a set of MuJoCo and VMAS continuous control tasks. Our experiments demonstrate that optimistic exploration significantly accelerates learning in environments with sparse rewards, action penalties, and difficult-to-explore regions. Furthermore, we provide insights into when optimism is beneficial and emphasize the critical role of model uncertainty in guiding exploration.
@@ 3 Problem Statement
for transition dynamics $f_{t}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and reward function $f_{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Our objective is to learn optimal control for this system within episodes of finite time horizon $T$. To control the agent, we learn a deterministic or stochastic policy $\pi$ that selects actions given the current state. For ease of notation, we write the policy deterministically as $\pi:\mathcal{S}\rightarrow\mathcal{A}$ so that $a_{t}=\pi(s_{t})$. We consider a specific transition dynamic $\hat{f}_{t}$ and reward function $\hat{f}_{r}$, denoted with a hat to indicate realizations that are drawn from random functions. The performance of the policy is measured by the sum of accumulated rewards during an episode,
@@ 5.1 Model Structure
where $\mu_{s}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and $\mu_{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ are the dynamics and reward mean functions and $K_{ss},K_{sr},K_{rr}$ represent covariance functions $(\mathcal{S}\times\mathcal{A})\times(\mathcal{S}\times\mathcal{A})\rightarrow\mathbb{R}$. Traditional GPs describe outputs as conditionally independent given the input and hence factorize across the state dimensions and the reward. However, we hypothesize that this assumption is unfounded in most RL tasks and furthermore antithetical to principled optimistic exploration. Therefore, we want to explicitly model the correlation structure between the dimensions of the predicted state, and more importantly, between the predicted state and its associated reward. While this correlation can naturally be included in the GP by parameterizing the full covariance structure across examples and output dimensions, this is computationally intractable due to cubic scaling with the size of the covariance matrix (Rasmussen & Williams, 2006). Instead, we adopt a linear model of coregionalization (Grzebyk & Wackernagel, 1994; Wackernagel, 2003) where the covariance between output dimensions is restricted to be linear combinations of the input covariance. This induces a Kronecker structured covariance function that allows for efficient factorization while still meaningfully directing optimistic exploration.
In most GP regression use cases, the mean function is a constant function. While this is a valid modeling assumption for simple data, the dynamics and reward structures we encounter in RL tasks vary in a non-constant manner across different regions of the input space. To better predict these complex transition-reward relationships, we use a GP model with a multi-layer perceptron (MLP) mean function, as in Iwata & Ghahramani (2017); Fortuin et al. (2019), and use the covariance function to learn the uncertainty, representing the residual around the MLP. This flexible model can capture intricate transitions while providing useful uncertainty estimates that can be leveraged in an optimistic fashion. The mean and covariance functions are learned jointly by performing approximate inference through optimizing a variational lower bound on the marginal likelihood, as done by Titsias (2009). This updates the posterior distribution $p(\hat{f}\mid\mathcal{D}_{\text{env}})$, which we use to obtain the posterior predictive distribution $p(s_{t+1},r_{t}\mid s_{t},a_{t})$. Given the current state $s_{t}$ and action $a_{t}$, this multivariate Gaussian distribution reflects the model’s predictions and uncertainty over the next state and reward.
@@ 5.2 Using Thompson Sampling in the GP Acquisition Function
Figure 1: Learning curves for all MuJoCo tasks were averaged over 10 seeds except for the Sparse Reacher task, which used 5 seeds. HOT-GP demonstrates equivalent or superior sample efficiency and performance for all tasks considered. The dashed line denotes SAC performance at convergence within 1,000,000 environment steps (3,000,000 for Half-Cheetah).
@@ 6 Experiments
We evaluate the performance of our proposed algorithm, HOT-GP, on continuous state-action control tasks by measuring the average sum of rewards accumulated during evaluation episodes. Since inaccurate reward estimates are likely initially, we increase $r_{\text{min}}$ linearly throughout training to become more optimistic as the model becomes more reliable. In our experiments, we consider three variations of HOT-GP where $r_{\text{min}}$ begins at 0.1 and increases to either 0.3, 0.5, or 0.7. We also consider the alternative Thompson sampling and greedy exploration strategies. Our Thompson sampling implementation follows the HOT-GP procedure but draws samples from the reward-conditioned distribution $p(s_{t+1}\mid s_{t},a_{t},r_{t}=\hat{r}_{t})$ rather than taking the expectation. For greedy exploitation, we predict the next state and reward as the mean from the model.
First, we compare these approaches on standard MuJoCo benchmark tasks (Todorov et al., 2012) and extended MuJoCo sparse maze tasks from the D4RL suite (Fu et al., 2020) against additional model-free and model-based methods. Then we evaluate these approaches on a practical robotics task using the VMAS simulator (Bettini et al., 2022) and provide an analysis of the individual components. We provide algorithmic descriptions of HOT-GP and the model-based comparison methods in Appendix A. Details about training are discussed in Appendix B, specific tasks are described in Appendix C, additional evaluations are given in Appendix D, and code for reproducing the experiments can be found at https://github.com/jbayrooti/hot_gp.
@@ 6.1 MuJoCo Benchmarks
Sparse Reacher. The Sparse Reacher task provides reward $r_{\text{dist}}(s)$ that is nonzero only when the agent is within a small threshold of the target location, thus demanding clever exploration strategies to reach the target as there is no dense feedback. As in Curi et al. (2020), we introduce an action penalty of the form $r(s,a)=r_{\text{dist}}(s)-\rho\cdot\text{cost}(a)$ with the aim of penalizing random and encouraging local exploration. This is described in greater detail in Appendix C. We evaluate HOT-GP and the other comparison methods on Sparse Reacher using four different action penalty weights: $\rho=0$, $\rho=0.1$, $\rho=0.3$, and $\rho=0.5$, and present the results in the bottom row of Figure 1 as well as Figure 7 in Appendix D. HOT-GP demonstrates significantly superior sample efficiency over MBPO and optimistic MBPO in each case. Also, the asymptotic performance of HOT-GP matches the SAC-best performance for each action penalty while achieving it in a fraction of the environment steps. When $\rho=0$, all HOT-GP variants perform robust optimistic exploration and successfully uncover the signal despite sparse rewards. When using the action penalty $\rho=0.1$, HOT-GP with $r_{\text{min}}$ from 0.1 to 0.3 performs marginally more poorly than with higher levels of optimism, suggesting that high levels of optimism may be required to overcome the prior against taking large actions and explore the reward signal $r_{\text{dist}}$. HOT-GP does this by relating the action-based penalty to the state to the overall reward through the joint model. The less efficient learning of optimistic MBPO highlights the critical importance of having a joint model of the reward and state distributions to facilitate efficient exploration. For $\rho=0.3$ and $\rho=0.5$, we observe that the cost term accounts for up to half of the overall reward. As a result, the HOT-GP variants explore much of the cost signal rather than $r_{\text{dist}}$, leading to similar performance in these settings.
Figure 2: Learning curves on sparse maze tasks averaged over 10 seeds on the U Maze and 5 seeds on the Medium Maze. HOT-GP achieves equivalent or superior sample efficiency and performance on both tasks. The dashed line denotes SAC performance after 2,000,000 environment steps.
Sparse Maze. In the Sparse Maze tasks, an agent must navigate through a maze to a randomized target location, receiving a reward of 1 when within a small threshold of the target location and zero in other cases. Similar to the Sparse Reacher task, efficiently solving such problems requires an effective exploration strategy due to the lack of dense feedback. We consider the simple U Maze and challenging Medium Maze environments, which are described in Appendix C. The results of HOT-GP and the comparison methods on these mazes are given in Figure 2. On the U Maze, HOT-GP attains equivalent sample efficiency and performance with the greedy and Thompson sampling strategies, while MBPO and optimistic MBPO are less stable. Greedy exploitation likely does well on this task because of the simple dynamics and reduced need for exploration. On the other hand, the Medium Maze presents a significantly more difficult exploration problem where greedy exploitation is insufficient to learn useful behavior. On the Medium Maze, HOT-GP with $r_{\text{min}}$ ranging from 0.1 to 0.7 delivers the strongest performance while other HOT-GP variants trail behind, showing that lower levels of optimism result in poorer performance. This disparity is likely due to the substantial sparsity in rewards and difficult maze layout, characterized by multiple dead ends, which necessitates more aggressive optimistic exploration. Thompson sampling initially matches the best HOT-GP performance, but later shows instability with a performance decline that may be due to stochastic next state sampling. In contrast, MBPO and optimistic MBPO learn less efficiently, further reinforcing HOT-GP’s advantage in jointly modeling reward and state distributions for effective exploration.
@@ 7 Conclusion
Future Work. Future work should study how the schedule on $r_{\text{min}}$ affects optimism and principled ways for setting this value throughout training, as we only considered a linear schedule. Another promising avenue would extend the planning horizon. We concentrated on learning 1-step lookahead reward-dynamics as HOT-GP can implicitly propagate uncertainty across the horizon by using pointwise uncertainty estimates from the model. Nevertheless, there may be benefits in explicitly propagating uncertainty to facilitate optimism over longer-term reward predictions (Seyde et al., 2021; Mehta et al., 2022; Hansen et al., 2024). Furthermore, our investigation indicates that the utility of optimistic exploration depends on the complexity of the task, although our experiments were limited to six tasks. Future research should explore the role of optimism across a broader range of problems, including in real-world scenarios with on-robot learning.
@@ Acknowledgments
J. Bayrooti is supported by a DeepMind scholarship. A. Prorok is supported in part by European Research Council (ERC) Project 949940 (gAIa). We also sincerely thank Markus Wulfmeier and Aidan Scannell for insightful conversations and valuable feedback, which significantly contributed to the development of this work.
@@ Appendix A Algorithm Details
In this section, we provide more detailed information about the algorithms we use for evaluation. Note that the code we used to implement these algorithms is available at https://github.com/jbayrooti/hot_gp.
@@ A.1 Model-Free Baselines
PPO. We use a standard version of PPO as introduced in Schulman et al. (2017).
@@ Appendix B Training
Model rollout steps $K$ 1 1 1 1 1
@@
Gaussian Process with Neural Network Mean Function. A standard Gaussian Process is defined by a mean function $m(x)$ and a covariance function $k(x,x^{\prime})$, where $x$ and $x^{\prime}$ are input vectors, and $m(x)$ is typically assumed to be a zero mean. In this work, we replace the traditional zero mean function with a neural network so that $m(x)=f_{\theta}(\mathbf{x})$, where $\theta$ are the parameters of the network. Given inputs $X=[x_{1},\dots,x_{n}]$ and outputs $y=[y_{1},\dots,y_{n}]$, the posterior distribution of this GP is
$p(y|X,y)=\mathcal{N}(m,K+\sigma^{2}I),$
where $m=[f_{\theta}(x_{1}),\dots,f_{\theta}(x_{n})]$ is the mean vector of neural network outputs for each input, K is the kernel matrix $K_{ij}=k(x_{i},x_{j})$, $\sigma^{2}$ is the noise variance, and $I$ is the identity matrix. To optimize this model, we follow the procedure in Iwata & Ghahramani (2017). This entails optimizing $\theta$ using the Mean Squared Error loss and then optimizing the kernel hyperparameters and noise variance $\sigma^{2}$ on the train dataset with transformed targets $y-m$, as described in the following section, on each iteration. For predicting on a new input $x_{*}$, the predictive posterior distribution is $p(y_{*}\mid x_{*},X,y)=\mathcal{N}(y_{*}\mid u_{*},\sigma_{*})$ with:
$\mu_{*}=m_{*}+K_{*}^{T}\left(K+\sigma^{2}I\right)^{-1}(y-m)$
$\sigma_{*}^{2}=K_{**}-K_{*}^{T}\left(K+\sigma^{2}I\right)^{-1}K_{*},$
where $m_{*}=f_{\theta}(x_{*})$, $\mathbf{K}_{*}=\left[k(x_{*},x_{1}),\dots,k(x_{*},x_{n})\right]^{T}$, $K_{**}=k(x_{*},x_{*})$. Thus, the neural network learns the non-zero mean function while the GP covariance function is modeled by the kernel as usual, thus allowing the model to quantify uncertainty around complex functions.
Reward-Dynamics Model Training. Our approach requires access to the full posterior distribution of the Gaussian process. To facilitate this in an efficient manner, we train the GP using a variational bound based on inducing points (Hensman et al., 2015) following the approach given in Iwata & Ghahramani (2017). This allows for efficient training using stochastic inference. To further reduce the computational cost of learning the model, we do not train on the whole dataset on each iteration and instead draw $1,000$ randomly sampled transitions from $\mathcal{D}_{\text{env}}$. This approach reduces the training cost significantly and provides an additional source of stochasticity that helps training. Using a Gaussian likelihood allows us to compute the predictive posterior in closed form, allowing us direct access to its mean and variance.
Hyperparameters. We train all algorithms using a set of hyperparameters that we found to be optimal through trial-and-error or were recommended. Table 1 describes the hyperparameter values specific to the MuJoCo and VMAS tasks we considered. We also report additional algorithm-specific hyperparameters used. In H-UCRL, we use the exploration-exploitation coefficient $\beta=0.01$ and number of samples $Z=5$. In MBPO and optimistic MBPO, we use an ensemble of size 7. For DDPG, we use an exploration noise $\sigma_{\text{explore}}$ from 1 to 0.1. For GP-based approaches, we use the Matern kernel as the covariance function and learn from 100 inducing points. For the MuJoCo tasks, we use a 4-layer MLP with 200 hidden units per layer and SiLU activation function. For the VMAS coverage task we use a 2-layer MLP with 200 hidden units per layer and Mish activation function. We use the Adam optimizer in all cases. For further information on our implementation, please see our open-sourced implementation at this repository:
https://github.com/jbayrooti/hot_gp.
Computation Time Comparison. We report the mean wall clock runtime and standard deviations for a single run of each algorithm on the MuJoCo tasks considered in Table 2. These results are averaged over 10 seeds. Our HOT-GP implementation incurs longer wall clock runtimes due to computational cost of GP training. However, recent research proposes promising techniques that can mitigate this overhead (Wenger et al., 2024; Chang et al., 2023). Also, note that our approach is not tied to GP models and can seamlessly generalize to alternative models such as Bayesian Neural Networks (BNNs), as the model is only needed to approximate the joint posterior distribution over next states and rewards.
Table 2: Computation times in hours across tasks and algorithms
Wall Clock Runtime MBPO Optimistic MBPO Greedy HOT-GP
Half-Cheetah 24.5 $\pm$ 3.5 26.7 $\pm$ 4.88 18.5 $\pm$ 3 28.89 $\pm$ 4.1
Reacher 1.89 $\pm$ 0.1 2.04 $\pm$ 0.2 2 $\pm$ 0.1 5.18 $\pm$ 0.2
Pusher 1.87 $\pm$ 0.2 2.05 $\pm$ 0.2 1.92 $\pm$ 0.4 6.53 $\pm$ 0.1
Sparse Reacher 3.76 $\pm$ 0.3 4.1 $\pm$ 0.4 3.94 $\pm$ 0.3 8.55 $\pm$ 1.5
U Maze 12.45 $\pm$ 1.1 12.43 $\pm$ 1.7 11.75 $\pm$ 1.6 21.08 $\pm$ 3.8